Find a 95% confidence interval for the current true proportion
More Prahces We want to estimate the proportion of Philadelphia families with children enrolled in preschool. In a random sample of 50 Philadelphia families with children of preschool age, 35 had children enrolled in preschool. Find a 95% confidence interval for the current true proportion of Philadelphia families with the children enrolled in preschool.
Answer #1

Solution:

Confidence Interval for population proportion(p)

Given,

n = 50 ....... Sample size

x = 35 .......no. of successes in the sample

Let hat p denotes the sample proportion.

  herefore  hat p = x/n = 35/50 = 0.70

Our aim is to construct 95% confidence interval.

herefore c = 0.95

hereforealpha = 1- c = 1- 0.95 = 0.05

herefore  alpha/2 = 0.05 slash 2 = 0.025 and 1- alpha /2 = 0.975

Search the probability 0.975 in the Z table and see corresponding z value

hereforeZ_{alpha/2} = 1.96 for the 95% Confidence interval

Now , confidence interval for population proportion(p) is given by:

p-Marqin.of.error<P<D+ Margin.of error

  hat p-Z_{alpha/2}*sqrthat p(1-hat p)/n   < p <hat p+Z_{alpha/2}*sqrthat p(1-hat p)/n

0.70 -1.96* уб.70(1-0.70),50 < p < 0.70 +1.96*уб.70(1-0.70),50

0.70 - 0.12702251768 <  p   < 0.70 + 0.12702251768

0.5730 < p <  0.8270

is the required 95% confidence interval for the population proportion.

Answer is (0.5730,0.8270)